RHIC cryo problems and shifts

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Sat Oct 20 2001 - 17:46:16 EDT

  • Next message: Apache: "Shift report 20011021 08:00-16:00"

    As you may have seen there are cryo-problems with RHIC. The latest estimate is operation will not
    be restored until Sunday morning
    
    It is therefore reasonable to suspend the evening and owl shift for this evening and night;
    The proper procedure for this is to
    -- a magnets to standby -> off 
    -- store keys in lock box
    -- Inform main control that no brahms shift person is available until Sunday Morning at 8.
    
    Please perform the regular evening shift check as well as a regular shift check first things Sunday morning
    
    
    Flemming
    
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    



    This archive was generated by hypermail 2b30 : Sat Oct 20 2001 - 17:43:57 EDT